THE TORSIONAL FUNDAMENTAL BAND OF METHYLFORMATE

M. TUDORIE, Service de Chimie Quantique et Photophysique, Université Libre de Bruxelles, CP 160/09, 50 avenue F.D. Roosevelt, B-1050 Brussels, Belgium; V. ILYUSHIN, Department of Microwave Radiospectrometry, Institute of Radio Astronomy of NASU, Chervonopraporna 4, 61002 Kharkov, Ukraine; \underline{J. VANDER AUWERA}, Service de Chimie Quantique et Photophysique, Université Libre de Bruxelles, CP 160/09, 50 avenue F.D. Roosevelt, B-1050 Brussels, Belgium; O. PIRALI, P. ROY, Ligne AILES – Synchrotron SOLEIL, L’Orme des Merisiers, F-91192 Gif-sur-Yvette, France; T. R. HUET, Laboratoire de Physique des Lasers, Atomes et Molécules, UMR CNRS 8523, Université Lille 1, 59655 Villeneuve d’Ascq Cedex, France.

Methylformate (HCOOCH_{3}) is one of the most important molecules in astrophysics, first observed in 1975.^{a} The rotational structure of its ground and first excited torsional states are well known from millimeter wave measurements.^{b} However, some of the torsional parameters are still not precisely determined because information on the torsional vibrational frequency v_{t} = 1 − 0 is missing. To overcome that problem, the far infrared spectrum of HCOOCH_{3} was recorded with a 150 m optical path in a White cell and a Bruker IFS 125 HR Fourier transform spectrometer at the AILES beamline of the synchrotron SOLEIL facility. The analysis of the very weak fundamental torsional band v_{t} = 1 − 0 observed around 130 cm^{−1} was carried out. It led to the first precise determination of the torsional barrier height and the dipole moment induced by the torsional motion.

This work is partly supported by the ”Programme National de Physico-Chimie du Milieu Interstellaire” (PCMI-CNRS) and by the contract ANR-BLAN-08-0054.

aR.D. Brown, J.G. Crofts, P.D. Godfrey, F.F. Gardner, B.J. Robinson, J.B. Whiteoak, Astrophys. J. 197 (1975) L29L31.

bSee V. Ilyushin, A. Kryvda, E. Alekseev, J. Mol. Spectrosc. 255 (2009) 32-38, and references therein.